OpenStudy (jamesreed30):

Express the required measurement in exact form a bicycle wheel with radius 26 inches rotates through an arc that measures 80° what is the length of the arc of the tire that touch the ground

OpenStudy (mhchen):

So you are given the radius, 26in. An angle, 80 deg. 1. Find the ratio of the angle measure to the entire angle, which is 80/360 2. Find the circumference, which is 2 * pi * radius Now given these, you can do a PROPORTION, right? \[\frac{anglè}{360} = \frac{arc}{circumference}\]

OpenStudy (mhchen):

\[\frac{80}{360} = \frac{arc}{2*\pi*26}\]

OpenStudy (jamesreed30):

Okay so what would I do next? @mhchen

OpenStudy (mhchen):

Can you solve the 'arc' in the proportion above?

OpenStudy (jamesreed30):

Divide 360 by 80?

OpenStudy (mhchen):

erm....like this:

OpenStudy (mhchen):

So you remember the proportion I made above right? If we multiple both sides by 2pi*radius, we get: \[\frac{80 * (2*\pi*26)}{360} = arc\]

OpenStudy (jamesreed30):

Okay I get that part &I also sorry Ima bit slow with this stuff this is all new to me